OpenStudy (anonymous):

Simplify (6x^18y^10)^3

OpenStudy (anonymous):

\[(6x ^{18}y ^{10})^{3}\]

OpenStudy (anonymous):

216y^30x^54 You multiply the degrees 3(18) and 10(3)

OpenStudy (anonymous):

wouldn't it be \[\frac{ x ^{6} }{ 216y ^{30}}\]

OpenStudy (unklerhaukus):

\[(ax^ny^m)^p=a^px^{np}y^{mp}\]

OpenStudy (anonymous):

Why would you be dividing?

OpenStudy (anonymous):

No clue that's just part of an example

OpenStudy (anonymous):

You just multiply each term by 3. 3(6), 3(18), 3(10)

OpenStudy (anonymous):

Your right ty

OpenStudy (anonymous):

Welcome. :)